Suppose you wanted to purchase a pair of boots. The original price is $59.99. They are on sale for 50% off the original price. You will then receive an extra 15% off the sale price.

a. First calculate the sales price by taking 50% off the original price and then take 15% off the sale price. What is the price of the boots after you apply both discounts?

b. How much did you save?

c. What is your overall percentage savings?
Answer #1

a)Price after taking 50% off the original price =59.99*(1-0.50)

price of the boots after you apply both discounts =59.99*(1-0.50)*(1-0.15)=$ 25.50

b)saved amount =59.99-25.50 =$ 34.49

c)Overall % saving =(saving/initial amount)*100 =(34.49/59.99)*100 =57.5 %